Help me with this problem, please. I want to understand

you will calculate the relative amount of photons passing through a slab. You can assume only all-or-nothing interactions occur The slab has a thickness of 2 em and is made of either aluminum or lead. Please use the following data when answering the question: μ = 0.27 cm-i and What percentage of the initial amount of photons will pass through the aluminum slab? Note: please do NOT type the percentage sign (%) in the answer box 시% 3.405 cm-1 And what percentage of the initial mount of photons will pass through the lead slab? Note: please do NOT type the percentage sign (96) in the answer box
Answer #1

Initial intensity of photons I_0

Intensity of photons after passing through a material of thickness x and absorption coefficient \mu isI=I_oe^{-\mu x}

Thickness x=2\,cm

For Aluminium, I=I_oe^{-\mu_{Al} x}

I=I_oe^{-0.27*2}=0.583I_0

\Rightarrow \frac{I}{I_0}*100=\frac{0.583I_0}{I_0}*100=58.3 \,\%

Percentage of initial amount of photons that pass through Aluminium is 58.3

---------------

For Lead , I=I_oe^{-\mu_{Pb} x}

I=I_oe^{-0.3.405*2}=0.0011\,I_0

\Rightarrow \frac{I}{I_0}*100=\frac{0.0011I_0}{I_0}*100=0.11 \,\%

Percentage of initial amount of photons that pass through Lead is 0.11

-----------------